OpenStudy (anonymous):

The local pizzeria sold 264 slices of pizza totaling $810.00. Cheese pizza costs $2.50 a slice and pepperoni pizza costs $3.25 a slice. How many slices of each pizza were sold? Let c represent cheese pizza and p represent pepperoni pizza. A. c = 64, p = 200 B. c = 72, p = 192 C. c = 132, p = 132 D. c = 200, p = 64

OpenStudy (anonymous):

@greenlegodude57 Can you help me?? Again??

OpenStudy (anonymous):

Yes, hold on

OpenStudy (anonymous):

This is SO easy, look. c = cheese pizza p = pepperoni pizza c = 2.50 a slice p = 3.25 a slice Option A: (2.50 x 64) + (3.25 x 200) = 160 + 650 = 810 So Option A is your answer!

OpenStudy (anonymous):

Its A. You wont get my explanation

OpenStudy (anonymous):

god I wish my stuff was that easy

OpenStudy (anonymous):

@chaser71 Ikr? *sigh*

OpenStudy (anonymous):

dont you wish all math was this easy? @greenlegodude57

OpenStudy (anonymous):

Yep, I finished Algebra I, but I still got a lot more to do..

OpenStudy (anonymous):

Yes, I wish it was!! =)

OpenStudy (anonymous):

Thank you guys for helping me!!! =)

OpenStudy (anonymous):

Who buys that much pizza. This can only happen in math. Lol